# Modular Compilation for Quantum Chiplet Architectures

### v1

@@ Abstract.
As quantum computing technology continues to mature, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of wildly varying levels of duration or fidelity. In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×. Additionally, owning to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
@@ 1. Introduction
We postulate that quantum computing is on a similar path. While improvements in superconducting quantum hardware have led to the debut of processors with $1000+$ qubits (ibm_condor), practical implementations of quantum computation will require millions of physical qubits (paler2019really). The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time, or number of gates on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of matter qubits when all control hardware is included, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC). The overhead for error correction, cooling, dilution, control systems, I/O lines, challenges associated with verification and adequate chip thermalization, and realistic resources such as lossy waveguides, limited qubit fabrication yields as qubit capacity grows (smith_2022_chiplets), and finite chip sizes, make the prospect of a single “monolithic” quantum processor very expensive or entirely unrealistic by current standards. These constraints force large-scale quantum systems to adopt a physically distributed architecture.
@@
Unfortunately, current compilation infrastructure is not capable of reasoning well about this quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution, as the ideal physical qubits to execute on change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in compilation complexity of $O(n^{2})$ for $n$ total qubits in a quantum processor.
@@ 2.2. Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets, and these links often don’t support universal operations (niu2023low). Thus, when algorithms like SABRE depend on the expectation of universality among qubit links, they may generate incompatible outputs when targeting modular architectures.
@@ 3. A Strawman Chiplet Compiler
(a) A physical inter-chiplet link exists between halo qubits $A_{0}$ and $B_{0}$. A physical intra-chiplet link exists between $B_{0}$ and $B_{1}$.
(b) A logical CNOT gate between chiplets $A$ and $B$ which is not implementable on hardware (indicated in red).
(c) To prepare for an inter-chiplet SWAP, the logical qubit on $B_{0}$ must be moved to a nearest neighbor, $B_{1}$ in this case, with a SWAP operation.
(d) The inter-chiplet SWAP between chiplets $A$ and $B$ completes transformation of the CNOT gate to an intra-chiplet gate.
(e) Once the CNOT gate has been successfully placed, the inserted SWAPs must be undone to not affect future gates in the circuit.
Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet halo edges.
It was challenging to establish an appropriate baseline for state-of-the-art quantum circuit compilation on chiplet-based quantum devices. Not only are modular quantum architectures relatively nascent, but compilation schemes for monolithic devices are largely incompatible with chiplet devices, as was discussed in greater detail in Section 2.
Thus, we took it upon ourselves to design a baseline compilation scheme for chiplet architectures, against which we compare our more-optimized design presented in Section 4. This “strawman” compiler augments existing compiler designs with the minimal modifications necessary to generate valid circuit outputs for chiplet architectures. Specifically, we insert a “peephole” compiler pass immediately after qubit routing, which corrects for the placement of two-qubit gates on inter-chiplet links as shown in Figure 1. This algorithm is inefficient, since it inserts a total of 4 SWAP gates — 2 inter-chiplet and 2 intra-chiplet — for every inter-chiplet gate. Unfortunately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
@@ 4. The Design of SEQC
We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure. The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
@@ 4.1. Stratification Stage
(a) A $3$-chiplet topology with $2$ qubits per chiplet ($6$ qubits in total).
(b) Mapping of a $6$-qubit quantum circuit to $3$ subcircuits. Gates represent arbitrary cross-subcircuit operations.
(c) Mapping of subcircuits to physical chiplets. Since chiplets $A$ and $C$ are not physically connected, certain gates (shown in red) necessitate inter-chiplet routing.
(d) Inter-chiplet routing is performed to address cross-chiplet gates that either don’t conform to the physical inter-chiplet topology or the basis gate set (aren’t SWAP gates).
Figure 2. Showcase of Inter-Chiplet Compilation in the Stratification Stage
In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 2(a), by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps, as shown in Figures 2(b)2(d). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
@@ 4.1.1. Qubit-to-Subcircuit Mapping
To perform qubit-to-subcircuit mapping (Figure 2(b)), we leverage a form of simulated annealing. Simulated annealing is a form of stochasic optimization, notably used in VLSI algorithms for place-and-route (annealing1989). We use simulated annealing to find a qubit assignment that minimizes the number of connections between subcircuits. As our chiplet allocation algorithm acts conservatively—performing the minimal number of changes necessary to make a circuit valid—it will produce better results if given a circuit with minimal connectivity between subcircuits. For our cost function, we count the number of inter-subcircuit gates. To permute the solution, we swap two random qubits in different subcircuits. As different simulated annealing runs are independent, running the algorithm using different seeds is embarrassingly parallel. We exploit this parallelism by running many trials at once, each in a different process with a different seed, then choosing the best result across all trials and seeds.
@@ 4.1.2. Chiplet Allocation
To perform chiplet-level allocation, shown in Figures 2(c) and 2(d), we extend and modify the SABRE algorithm for initial mapping (sabre_swap) with application-specific layout and routing passes. For layout, we initialize each iterative allocation cycle with a random placement of logical subcircuits into physical chiplets. For routing, we developed a new heuristic for deciding between candidate SWAPs.
By default, SABRE attempts to place two-qubit gates on a valid edge in the graph of a device’s hardware topology, i.e, the physical distance between the qubits of a gate should be $1$. For inter-chiplet routing, we do not wish to factor intra-qubit topology into our distance calculation. Furthermore, we hope to move all gates to be within the same chiplet, i.e., the distance between qubits should be 00 instead of $1$. Thus, when selecting between potential SWAPs, we categorize and prioritize SWAPs in a three-tier approach:
  1. (1)
    Symbiotic SWAPs: A SWAP between two multi-chiplet gates is considered symbiotic if it reduces the distance of both gates, i.e., each gate benefits from the SWAP. In theory, we should expect symbiotic SWAPs to be the most efficient, and thus highest-priority, SWAP possible.
  2. (2)
    Commensalistic SWAPs: When a SWAP can benefit one multi-chiplet gate without harming any other, we consider it to be commensalistic. Such behavior occurs most commonly on idle qubits in the quantum circuit, although it can occur between two multi-chiplet gates.
  3. (3)
    Parasitic SWAPs: The remaining category of SWAPs to consider benefit one gate while harming another. Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances. For example, clear one gate may free physical qubits to help route others.
With these changes, our algorithm is able to correctly and efficiently place and route circuits for any valid device topology.

### v2

@@ Abstract.
As quantum computing technology continues to mature, industry is adopting modular quantum architectures to keep quantum scaling on the projected path and meet performance targets. However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Contemporary compilation methods are not well-suited to chiplet architectures — in particular, existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set. Furthermore, existing methods of logical-to-physical qubit placement, swap insertion (routing), unitary synthesis, and/or optimization are typically not designed for qubit links of significantly varying levels of duration or fidelity. In this work, we propose SEQC, a complete and parallelized compilation pipeline optimized for chiplet-based quantum computers, including several novel methods for qubit placement, qubit routing, and circuit optimization. SEQC attains up to a 36% increase in circuit fidelity, accompanied by execution time improvements of up to 1.92×1.92$\times$1.92 ×. Additionally, owing to its ability to parallelize compilation, SEQC achieves consistent solve time improvements of 24×2-4$\times$2 - 4 × over a chiplet-aware Qiskit baseline.
@@ 1. Introduction
We postulate that quantum computing is on a similar path. While improvements in superconducting quantum hardware have led to the debut of processors with $1000+$ qubits (ibm_condor), practical implementations of quantum computation will require millions of physical qubits (paler2019really). The number of qubits required by a quantum algorithm, the depth of the algorithm (i.e., execution time or gate count on the critical path), the auxiliary and syndrome qubits required to reach sufficiently low error rates, and the size of qubits with all control hardware, all suggest that many more qubits are required than are likely to fit in a single die (vanMeter_2010_DQC). The overhead for error correction, cooling, dilution, control systems, I/O lines, challenges associated with verification and adequate chip thermalization, and realistic resources such as lossy waveguides, limited qubit fabrication yields as qubit capacity grows (smith_2022_chiplets), and finite chip sizes, make the prospect of a single “monolithic” quantum processor very expensive or entirely unrealistic by current standards. These constraints force large-scale quantum systems to adopt a physically distributed architecture.
@@
Unfortunately, current compilation infrastructure is not capable of reasoning well about this quantum interconnect heterogeneity. The mapping of a quantum program’s logical quantum gates into the native gates supported by the underlying quantum processor, and the mapping of logical qubits to the physical qubits of the quantum hardware, largely determine the number and type of native quantum gates required for the computation, the circuit depth, and its execution time—long programs may not complete successfully as qubits decohere. Different mappings can result in vastly different compiled quantum circuits, with diverse characteristics along all these axes. These details of efficiency can make or break an algorithm in today’s noisy intermediate-scale quantum systems (NISQ) era, so quantum compilers aggressively optimize for all of them. To make matters more challenging, each qubit, coupler, and gate have diverse error profiles that are highly variable both spatially and temporally (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Unlike classical compilation that is done only once for a given architecture, quantum programs must be recompiled every time before execution, as the ideal physical qubits to execute on change between runs. So, quantum compilers today are faced with the daunting task of optimizing quantum programs across multiple dimensions with often conflicting demands, in a continuously changing environment. Coupled with the hardness of synthesizing unitaries (which is exponential in the number of qubits) and the need for quick and frequent compilation, modern quantum compilers have no option other than to rely on heuristics to perform the task. These heuristics result in a compilation complexity that we experimentally observed to be $O(n^{2})$ for $n$ total qubits in a quantum processor, as shown in Figure 8.
@@ 2. Background and Motivation
(a) A chiplet quantum architecture
(b) Input circuit from SABRE
(c) Insertion of perfect-shuffle
(d) Insertion of return SWAPs
Figure 1. Peephole correction for incompatible (non-SWAP) gates placed on inter-chiplet edges. Gates that are unimplementable on hardware are shown in red.
@@ 2.2. Basis Translation
Basis translation converts the gates of an input quantum circuit into gates that are physically supported on hardware. Monolithic quantum architectures typically support a so-called universal basis gate set, where any unitary quantum operation can be constructed using solely the gates in the set (mike_and_ike). Moreover, it has been shown that basis translation can be performed efficiently given the target gate set is universal, and the quantum circuit contains a fixed, finite number of qubits (Solovay–Kitaev). Unfortunately, modular quantum architectures depend upon specialized links to connect otherwise-independent chiplets. In the general case, the cross-chiplet links can only be used for data movement (such as SWAPs) and not computation, meaning these links may not support universal operations (niu2023low). Thus, compilers that depend on the expectation of universality among qubit links (qiskit_basis_translation_errors) and do not expect these heterogenous, non-universal basis gate sets can generate incompatible outputs when targeting modular architectures.
@@ 3. A Strawman Chiplet Compiler
Figure 2. High-Level Diagram of SEQC. Elements in red reflect modifications from existing compiler designs.
As discussed previously in Section 2, our model of modular quantum architectures includes the more general case where inter-chiplet links are can be non-universal. This is representative of hardware manufacturing constraints (niu2023low), where only a restricted set of operations, e.g., state transfers or SWAPs, may be supported over long-distance couplers (inter-chiplet links). However, many compiler designs (including designs for modular architectures, see Section 7) experience errors or crashes when presented with non-universal, heterogenous, basis gate sets. For example, the Qiskit compiler (qiskit2024) is explicit in its compatibility for only universal basis gate sets (qiskit_basis_translation_errors). While it is possible to create a Qiskit Backend that supports heterogeneous basis gate sets, the compiler itself does not take them into account. Specifically, the Qiskit routing stage (zou2024lightsabrelightweightenhancedsabre) (an implementation of SABRE) will freely place gates on incompatible physical links, and there are no methods in the compiler to prevent or fix consequent errors. For the purposes of this work, however, it was critical to establish a baseline representative of the current state of quantum compilation, against which we can compare our more-optimized design (see Section 4). Thus, we created a “Strawman” baseline compiler, built on top of the Qiskit compiler (qiskit2024), with the minimal modifications necessary to generate valid outputs for chiplet architectures.
To accomplish this, we leverage a “peephole” compilation pass, which is commonly used in classical compilation to perform optimization (dragon_book). Peephole passes operate within a nominal peephole window and substitute its contents with a logically-equivalent replacement. Notable examples in the context of quantum compilation include basis translation and elimination of self-adjoint gates. This design strategy facilitates modular compilation, not only in the sense that it can be easily parallelized, but also because peephole passes can be added between other compilation passes without breaking compatibility. These qualities make a peephole pass well-suited to the Strawman baseline, since we do not want to make extensive modifications to any of Qiskit’s built-in compilation stages or passes, nor introduce a significant performance overhead. The peephole pass used in the Strawman compiler is inserted between the routing and translation compilation stages in Qiskit. As shown in Figure 1, it operates on the circuit generated by SABRE (zou2024lightsabrelightweightenhancedsabre), where it detects when non-SWAP gates are placed on non-universal inter-chiplet links and employs a cyclic qubit rotation, i.e., a perfect shuffle permutation, to move the aforementioned gates to nearby intra-chiplet links ( see Figure 1(c)). Overall, for every inter-chiplet gate in the original circuit generated by SABRE, the peephole pass inserts 4 additional SWAP gates—2 inter-chiplet and 2 intra-chiplet. Ultimately, this degree of overhead is inevitable when the stock, unmodified compiler is ignorant of chiplet-specific device constraints.
To emphasize the rigor of the Strawman, it is worth discussing the designs that were ultimately discarded. For example, consider a hypothetical alternative design that skips the final peephole step shown in Figure 1(d). This would appear to save $2$ SWAPs, thus halving the cost per inter-chiplet gate. However, the hypothetical pass would also no longer be a true peephole, since it introduces a net qubit permutation that propagates to downstream gates in the circuit. From the perspective of modular compilation, this results in serial dependencies that prevent parallel execution of the pass. From a routing perspective, the resultant qubit permutations disrupt the solution generated by SABRE, leading to destructive routing errors that can only be resolved by an entirely new routing pass. Overall, such a hypothetical pass results in substantially worse output circuits compared to the peephole method, which we confirmed through testing. Other potential improvements on the peephole design that avoid the placement of return SWAPs would involve extensive modifications to Qiskit’s compilation stages or its underlying passes. In Section 4, we will explore such a modification of SABRE that accounts for the only-the-fly changes in qubit layout associated with inter-chiplet routing as an element of our design, in addition to numerous other components to optimize compilation for modular quantum architectures.
@@ 4. The Design of SEQC
We incorporate the hierarchical elements of chiplet architectures into our compiler design by adopting a two-stage procedure, as shown in Figure 2. The first, stratification stage, represents a one-time overhead to elevate a quantum circuit to the chiplet level. The second, elaboration stage, represents a recurring cost to complete the compilation of a quantum circuit given the most up-to-date backend properties.
@@ 4.1. Stratification Stage
(a) A chiplet topology
(b) Qubit-to-Subcircuit Mapping
(c) Subcircuit-to-Chiplet Routing
(d) Inter-Chiplet Routing
Figure 3. Step-by-step example of inter-chiplet compilation in the Stratification stage.
In this stage, we map a quantum circuit to the inter-chiplet device topology, e.g., the structure in Figure 3(a), by stratifying it into smaller, interconnected subcircuits. This can be accomplished with the following intermediate steps, as shown in Figures 3(b)3(d). First, qubit-to-subcircuit mapping designates which logical qubits belong to which logical subcircuits. Then, as part of chiplet allocation, we define a mapping of logical subcircuits to physical chiplets on the target device and route inter-chiplet gates to conform with constraints of physical-chiplet connectivity.
@@ 4.1.1. Qubit-to-Subcircuit Mapping
Algorithm 1 Qubit-to-Subcircuit Mapping
[…]
2:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
3:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
4:Increment $C$ if $s_{0}\neq s_{1}$EndFor
[…]
7:$s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
8:$q_{0}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}])$
9:$s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
10:$q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{1}])$LCommentSwap $q_{0}$ and $q_{1}$
11:$L^{\prime}[s_{0}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\}$
12:$L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$EndFor
[…]
14:$L\leftarrow$ Initial Layout $\triangleright$ Current Solution
15:$R\leftarrow L$triangleright Best Solution
[…]
17:$D\leftarrow\textsc{Ceil}(T\times\alpha)$triangleright Number of changes to make
18:$L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
19:$\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
20:$k\leftarrow\textsc{Rand}(0,1)$triangleright Random number in $[0,1)\subset\mathbb{R}$
21:$L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
22:$R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
[…]
24:\$R$EndProcedure
[…]
Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits are grouped into subcircuits, as shown in Figure 3(b). An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution. To perform this task, we leverage simulated annealing, an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989).
Simulated annealing is an iterative method to optimize a state against a given cost function. In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost. However, the algorithm may randomly also accept worse solutions to escape from local minima wells. The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial. Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result. This improves the chances of finding a (near-)optimal solution.
Our method is shown in Algorithm 1. We start with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined on line 1 as the number of inter-chiplet gates given a certain layout. By minimizing the number of inter-chiplet gates, we implicitly reduce the inter-chiplet connectivity. The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits. In doing so, we hope to move strongly interacting qubits close-together, minimizing the number of cross-chiplet gates. As input, our algorithm takes the input circuit, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling modifier $\beta$. The last three are tuning parameters that do not change across runs.
@@ 4.1.2. Chiplet Allocation
Algorithm 2 Inter-Chiplet Routing
[…]
8:$V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
[…]
12:$L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$
[…]
14:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
15:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
[…]
18:$D\leftarrow\textsc{CircuitToDAG}(Circ)$
19:$E\leftarrow$ All-pairs distances from chiplet topology $T$
20:$Q\leftarrow\textsc{FirstLayer}(DAG)$
21:$R\leftarrow$ empty list
[…]
23:$\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$triangleright SWAP Candidates
24:$S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
25:$\textsc{Swap}(L,S)$
26:$\textsc{Step}(Q,R,L,D)$EndWhile
27:$D^{\prime}\leftarrow$ Empty DAG
28:Append $g$ to $D^{\prime}$ for each $g\in{R}$triangleright $R$ is ordered
[…]
Chiplet allocation describes the process of transforming a stratified circuit to be compatible with the inter-chiplet topology of a modular device. In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing. Accordingly, subcircuit-to-chiplet mapping, shown in Figure 3(c), defines an initial placement of logical subcircuits to physical chiplets. Subsequently, inter-chiplet routing, shown in Figure 3(d), inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
In our design, we do not explictly derive a subcircuit-to-chiplet mapping. Rather, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, thereby using iterations of inter-chiplet routing with randomized permutations to inform a subcircuit-to-chiplet mapping. Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit. For an single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph. From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain.
However, in the context of a entire stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave otherwise efficiently reorder execution of SWAP chains to reduce the depth and/or total number of inter-chiplet SWAPs. Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time. Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic. For our inter-chiplet routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 2. An inter-chiplet SWAP is considered symbiotic if it directly affects two inter-chiplet gates in a “beneficial” manner. In other words, each qubit of a symbiotic SWAP corresponds to a different inter-chiplet gate, where the insertion of the SWAP reduces the chiplet distance score of both gates. In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-chiplet routing cost heuristic. Next, commensalistic SWAPs are those that do not “harm,” or increase that chiplet distance, of any directly impacted gates. These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates. In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commercialistic behavior. Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed. Since this variety of SWAP is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock. For example, clearing one gate may free physical qubits to help route others.